# FAST drift scan survey for Hi intensity mapping: II. stacking-based primary beam construction for FAST L-band 19 feeds at $1.4$ GHz

### v1

@@ Abstract
Neutral hydrogen (Hi) intensity mapping (IM) presents great promise for future cosmological large-scale structure surveys. However, a major challenge for Hi IM cosmological studies is to accurately subtract the foreground contamination. An accurate beam model is crucial for improving foreground subtraction accuracy. In this work, we develop a stacking-based beam reconstruction method utilizing the radio continuum point sources within the drift-scan field. Based on the Five-hundred-meter Aperture Spherical radio Telescope (FAST), we employ two series of drift-scan survey data and merge the measurements to construct the beam patterns of the FAST L-band 19 feeds. To model the beams, we utilize the Zernike polynomial (ZP), which effectively captures asymmetric features of the main beam and the different side lobes. Due to the symmetric location of the beams, the main features of the beams are closely related to the distance from the center of the feed array, e.g., as the distance increases, side lobes become more pronounced. This modeling pipeline leverages the stable drift-scan data to extract beam patterns exclude the reflector’s changing effects, and provides a more accurate measurement beam and a more precise model beam for FAST Hi IM cosmology surveys.
@@ 1 Introduction
Neutral hydrogen intensity mapping (Hi IM), which is to measure the total Hi intensity of many galaxies within large voxels (e.g. Battye et al., 2004; McQuinn et al., 2006), has been proposed as a novel method for cosmic large-scale structure surveys in the epoch of post-reionization (Chang et al., 2008; Loeb & Wyithe, 2008; Mao et al., 2008; Pritchard & Loeb, 2008; Wyithe & Loeb, 2008; Wyithe et al., 2008; Peterson et al., 2009; Bagla et al., 2010; Seo et al., 2010; Lidz et al., 2011; Ansari et al., 2012; Pritchard & Loeb, 2012; Battye et al., 2013). It can be rapidly conducted and expanded to cover substantial survey areas, making it well-suited for cosmological surveys (Xu et al., 2015; Jin et al., 2020; Zhang et al., 2021; Jin et al., 2021; Wu & Zhang, 2022; Wu et al., 2023b, a; Zhang et al., 2023; Pan et al., 2024).
The Hi IM technique was investigated by measuring the cross-correlation function between an Hi IM survey and an optical galaxy survey. Chang et al. (2010) for the first time reports the cross-correlation function detection using the Hi IM carried out with Green Bank Telescope (GBT) and an optical galaxy survey. Later, several detections of the cross-correlation power spectrum between an Hi IM survey and an optical galaxy survey were reported (Masui et al., 2013; Anderson et al., 2018; Wolz et al., 2017, 2022; Amiri et al., 2023). Recently, the MeerKAT Hi IM survey (Bull et al., 2015; Cunnington et al., 2021; Li et al., 2021; Wang et al., 2021; Paul et al., 2021; Chen et al., 2023) reported the cross-correlation power spectrum detection with the optical galaxy survey (Cunnington et al., 2023). Meanwhile, the Hi IM auto power spectrum on Mpc scales is detected using the MeerKAT interferometric observations (Paul et al., 2023). However, the Hi IM auto power spectrum on large scales remains undetected (Switzer et al., 2013). Besides, several Hi IM experiments are targeting the post-reionization epoch, with some currently collecting data, including projects like the Tianlai project  [^1] https://tianlai.bao.ac.cn (Chen, 2012; Li et al., 2020; Wu et al., 2021; Perdereau et al., 2022; Sun et al., 2022) and the Canadian Hydrogen Intensity Mapping Experiment  (CHIME  [^2] https://chime-experiment.ca/en; Bandura et al., 2014), while others are in the construction phase, such as the Baryonic Acoustic Oscillations from Integrated Neutral Gas Observations  (BINGO  [^3] https://bingotelescope.org; Battye et al., 2013) and the Hydrogen Intensity and Real-time Analysis eXperiment  (HIRAX  [^4] https://hirax.ukzn.ac.za; Newburgh et al., 2016). The Hi IM technique is also proposed as the key cosmology project with Square Kilometre Array  (SKA  [^5] https://www.skao.int; Santos et al., 2015; Square Kilometre Array Cosmology Science Working Group et al., 2020). With the Five-hundred-meter Aperture Spherical radio Telescope, (FAST; Nan et al., 2011; Li & Pan, 2016) Hi IM survey shows considerable potential for cosmological studies (Li & Ma, 2017; Hu et al., 2020).
@@
The primary beam effect refers to two important aspects of the antenna beam pattern. The first is the main lobe shape and its frequency dependence; the second is the side-lobe beam pattern level and its frequency-dependent variation. Both can weaken the smooth foreground spectrum assumption, resulting in a failed foreground subtraction (Matshawule et al., 2021; Spinelli et al., 2022). To address such primary beam effects, the observed Hi maps are usually degraded to a common beam size, i.e. the largest beam size within the observation bandwidth, with an assumption of simple Gaussian primary beam shape (Masui et al., 2013; Anderson et al., 2018; Cunnington et al., 2023). This approach can reduce the primary beam effect to a subdominant level while also significantly suppressing the signal-to-noise ratio. Recently, a deep-learning-based strategy has been developed with simulated data for reducing the Hi IM survey systematic effect, which includes both the primary beam effect (Ni et al., 2022) and the polarization leakage effect (Gao et al., 2023). To apply such a strategy to the realistic observation data, an accurate beam model is required for training the deep learning network.
The beam pattern can be properly measured via radio holographic measurements with an interferometer (Iheanetu et al., 2019; Asad et al., 2021; Amiri et al., 2024), or by scanning a bright celestial calibrator for a single dish telescope (e.g. Jiang et al., 2020). Besides, the Tianlai experiment has measured the primary beam via unmanned aerial vehicle (Zhang et al., 2021) and CHIME has measured the primary beam via the sun transition (Amiri et al., 2022). The FAST beam pattern combines the response pattern of the coaxial feed and the reflector. The FAST L-band observation adopts a 19-feed array mounted in the focus cabin suspended on the cables about the reflector. The reflector consists of about 4 40044004\,4004 400 active reflector panels, which form a $300$ m paraboloid at each pointing direction. Such a complicated reflector and feed arrangement heavily increases beam pattern instabilities, resulting in a direction-dependent beam pattern (Dong & Han, 2013). In this work, we developed a stacking-based beam pattern construction method employing Hi IM drift scan observation data itself. Because the beam shape is measured in the same direction as the Hi IM survey observation, we may minimize the direction dependency impact and achieve the beam pattern that is close to the actual case during the observation.
@@ 2 Data and preprocessing
Table 1: The observational data used in this work. In these observations, the low-level noise diode is adopted and the feed array is rotated by $23.4^{\prime}$ during each observation.
@@
    The FATHOMER observations are carried out for $4$ hours drift-scan every nights with $1$ second integration time.
@@
    The CRAFTS observations are carried out for $5$ hours drift-scan every nights with $0.2$ second integration time.
@@
    The field center, encoded with ‘HHMM+ddmm’, indicates the center of each drift scan survey field, where ‘HHMM’ represents the R.A. and ‘+ddmm’ represents the decl., respectively.
@@ 2.1 FATHOMER
We employ the drift scan observation data from the FAst neuTral HydrOgen intensity Mapping ExpeRiment (FATHOMER; Li et al., 2023). In this analysis, $18\times 4$ hr drift scan observation data with integration time of $1$ s per timestamp are used, which targets the area that covers the R.A. range from 9 to 13 hr, overlapping with the Northern Galactic Cap (NCP) area of the Sloan Digital Sky Survey (SDSS; Reid et al., 2016). The drift scan time-ordered data are preprocessed via the analysis pipeline developed in Li et al. (2023). Following the pipeline, the full frequency band is first split into three sub-bands, i.e. the low-frequency band 10501050105010501150115011501150 MHz, the mid-frequency band 11501150115011501250125012501250 MHz, and the high-frequency band 12501250125012501450145014501450 MHz. In this work, we use the high-frequency band data as it is less contaminated by RFI. The high-frequency band also covers the continuum spectrum range of the celestial point source catalog, which is used as a sky model for beam pattern construction. Extending the lower frequency band requires a better knowledge of the source spectrum model. The selected data are then RFI flagged and flux calibrated before proceeding to the analysis of this work.
@@ 2.2 CRAFTS
The Commensal Radio Astronomy FasT Survey (CRAFTS; Li et al., 2018) is designed to cover the FAST sky between decl. of $-14^{\circ}$ and $+66^{\circ}$ and proposed to carry out multiple surveys simultaneously in a single drift scan observation, including Galactic Hi, extra-galactic Hi, pulsar search, and transient surveys. This analysis uses $14\times 5$ hr CRAFTS drift scan observation data, which targets the area that covers the R.A. range from $1$ to $6$ hr and the decl. range from $27^{\circ}15^{\prime}$ to $32^{\circ}19^{\prime}$. CRAFTS drift scan observation data have the high-cadence noise diode signal injection and the noise diode temperature was pre-calibrated (Jiang et al., 2020). The integration time for CRAFTS is $0.2$ s per timestamp. We reduce the integration time to $1$ s after the calibration processing. We adopt the same frequency partition strategy and time-ordered data analysis pipeline as the FATHOMS analysis. All the FAST drift scan data used in this work are summarized in Table 1.
@@ 2.3 NVSS
The FATHOMER and CRAFTS survey fields are both fully covered by the NVSS field. Both the FATHOMER and CRAFTS surveys adopt the drift scan observation mode, which is scanning at a constant decl. within an R.A. range. During the drift scans, the 19-feed array is rotated by $23.4^{\circ}$ to obtain the optimized field coverage (Li et al., 2018). Therefore, the same sources might be observed multiple times by different feeds. In addition, we also selected the sources with their decl. within $10^{\prime}$ to the scanning decl. to recover a full pattern beyond the size of the main beam. We choose a total area of $\sim 138\,{\rm\deg}^{2}$ and $\sim 385\,{\rm\deg^{2}}$ NVSS field for FATHOMER and CRAFTS, respectively. The total numbers of sources within the survey areas are 7 46374637\ 4637 463 and 21 0152101521\ 01521 015 for FATHOMER and CRAFTS, respectively. However, not all the sources are used in the beam pattern construction. Some of the sources are dropped because they are either poorly observed or highly contaminated by RFI or nearby sources. The detailed source selection criteria are presented in Section 3.3.
@@ 3.1 Beam stacking
where $\bm{d}$ and $\bm{n}$ are the time-ordered data vector and the noise vector, with the vector element representing the flux measurements and the corresponding noise at each time stamp, respectively. The multiplication of $\bm{M}\bm{b}$ represents the convolution of the sky model and the beam kernel. $\bm{b}=\{\bm{B}_{0}^{\rm T},\bm{B}_{1}^{\rm T},\cdots,\bm{B}_{q}^{\rm T}\}^{\rm T}$ is a column vector by stacking the columns of beam matrix $\bm{B}$ with size $(q,q)$, where $\bm{B}_{i}$ represents the $i$-th column of the beam matrix. Matrix $\bm{M}$ is a special matrix of size $(p,q^{2})$, where $p$ is the number of pixels within the survey area. the matrix $\bm{M}$ captures the convolutional spatial relationships between the pixels of the sky and the beam kernel matrix $\bm{B}$, i.e. each row of the matrix $\bm{M}$ is contracted by stacking the pixels covered by the beam kernel at each pointing direction. The matrix $\bm{P}$ is the projection matrix that maps the beam-convolved sky pixels to the time stamps.
@@
where $\hat{b}_{i,\cdots,j}$ represents a vector composed with the $i$-th to $j$-th elements of $\hat{\bm{b}}$.
In this work, the beam pattern is represented with a $100\times 100$ matrix, covering $20^{\prime}$ in both dimensions, which is large enough to include the first few side lobes.
@@ 3.2 The sky model
$m(\alpha,\beta)=\sum_{i=1}^{N_{\rm S}}S_{i}\delta_{\alpha\alpha_{i},\beta\beta_{i}},$ (5)
where $\delta_{\alpha\alpha_{i},\beta\beta_{i}}$ is the Kronecker delta function, $m(\alpha,\beta)$ is the flux density of the map, $(\alpha,\beta)$ represent the coordinate of the map pixel, i.e. R.A., and decl., $(\alpha_{i},\beta_{i})$ represents the coordinate of the $i$-th sources with flux density of $S_{i}$, and $N_{i}$ is the total number of sources in the chosen area. We use all the sources with flux density over $20\,{\rm mJy}$ to construct the sky model.

### v2

# FAST drift scan survey for Hi intensity mapping: II. stacking-based beam construction of the 19-feed array at $1.4$ GHz

@@ Abstract
Neutral hydrogen (Hi) intensity mapping (IM) presents great promise for future cosmological large-scale structure surveys. However, a major challenge for Hi IM cosmological studies is to accurately subtract the foreground contamination. An accurate beam model is crucial for improving the quality of foreground subtraction. In this work, we develop a stacking-based beam reconstruction method utilizing the radio continuum point sources within the drift-scan field. Based on the Five-hundred-meter Aperture Spherical radio Telescope (FAST), we employ two sets of drift-scan survey data and merge the measurements to construct the beam patterns of the 19 FAST L-band feeds. To model the beams, we utilize the Zernike polynomial (ZP), which effectively captures asymmetric features of the main beam and the different side lobes. Due to the symmetric location of the beams, the main features of the beams are closely related to the distance from the center of the feed array, e.g., as the distance increases, side lobes become more pronounced. This modeling pipeline leverages the stable drift-scan data to extract beam patterns while accounting for and excluding the reflector’s changing effects. It provides a more accurate measurement beam and a more precise model beam for FAST Hi IM cosmology surveys.
@@ 1 Introduction
Neutral hydrogen intensity mapping (Hi IM), which is to measure the total Hi intensity of many galaxies within large voxels (e.g. Battye et al., 2004; McQuinn et al., 2006), has been proposed as a novel method for cosmic large-scale structure surveys in the epoch of post-reionization (Chang et al., 2008; Loeb & Wyithe, 2008; Mao et al., 2008; Pritchard & Loeb, 2008; Wyithe & Loeb, 2008; Wyithe et al., 2008; Peterson et al., 2009; Bagla et al., 2010; Seo et al., 2010; Lidz et al., 2011; Ansari et al., 2012; Pritchard & Loeb, 2012; Battye et al., 2013). It can be rapidly conducted and expanded to cover substantial survey areas, making it well-suited for cosmological surveys (Xu et al., 2015; Jin et al., 2020; Zhang et al., 2021; Jin et al., 2021; Wu & Zhang, 2022; Wu et al., 2023b, a; Zhang et al., 2023; Pan et al., 2025).
The Hi IM technique was investigated by measuring the cross-correlation function between an Hi IM survey and an optical galaxy survey. Chang et al. (2010) first reported a cross-correlation detection using the Hi IM obtained with the Green Bank Telescope (GBT). Subsequent studies have also observed cross-correlation power spectra using various telescopes (Masui et al., 2013; Anderson et al., 2018; Wolz et al., 2017, 2022; Amiri et al., 2023). Recently, the MeerKAT Hi IM survey (Bull et al., 2015; Cunnington et al., 2021; Li et al., 2021; Wang et al., 2021; Paul et al., 2021; Chen et al., 2023) reported the cross-correlation power spectrum detection with the optical galaxy survey (Cunnington et al., 2023; MeerKLASS Collaboration et al., 2025; Carucci et al., 2024). Meanwhile, the Hi IM auto power spectrum on Mpc scales is detected using the MeerKAT interferometric observations (Paul et al., 2023). However, the Hi IM auto power spectrum on large scales remains undetected (Switzer et al., 2013). Besides, several Hi IM experiments are targeting the post-reionization epoch, with some currently collecting data, including projects like the Tianlai project  [^1] https://tianlai.bao.ac.cn (Chen, 2012; Li et al., 2020; Wu et al., 2021; Perdereau et al., 2022; Sun et al., 2022) and the Canadian Hydrogen Intensity Mapping Experiment  (CHIME  [^2] https://chime-experiment.ca/en; Bandura et al., 2014), while others are in the construction phase, such as the Baryonic Acoustic Oscillations from Integrated Neutral Gas Observations  (BINGO  [^3] https://bingotelescope.org; Battye et al., 2013) and the Hydrogen Intensity and Real-time Analysis eXperiment  (HIRAX  [^4] https://hirax.ukzn.ac.za; Newburgh et al., 2016). The Hi IM technique is also proposed as the key cosmology project with Square Kilometre Array  (SKA  [^5] https://www.skao.int; Santos et al., 2015; Square Kilometre Array Cosmology Science Working Group et al., 2020). With the Five-hundred-meter Aperture Spherical radio Telescope, (FAST; Nan et al., 2011; Li & Pan, 2016) Hi IM survey shows considerable potential for cosmological studies (Li & Ma, 2017; Hu et al., 2020).
@@
The primary beam effect refers to two important aspects of the antenna beam pattern. The first is the main lobe shape and its frequency dependence; the second is the side-lobe beam pattern level and its frequency-dependent variation. Both can weaken the smooth foreground spectrum assumption, resulting in a failed foreground subtraction (Matshawule et al., 2021; Spinelli et al., 2022). To address such primary beam effects, the observed Hi maps are usually degraded to a common beam size, i.e. the largest beam size within the observation bandwidth, with an assumption of simple Gaussian primary beam shape (Masui et al., 2013; Anderson et al., 2018; Cunnington et al., 2023). This approach can reduce the primary beam effect to a subdominant level while also significantly increasing the signal-to-noise ratio. Recently, a deep-learning-based strategy has been developed with simulated data for reducing the Hi IM survey systematic effect, which includes both the primary beam effect (Ni et al., 2022) and the polarization leakage effect (Gao et al., 2023). To apply such a strategy to the realistic observation data, an accurate beam model is required for training the deep learning network.
The beam pattern can be properly measured via radio holographic measurements with an interferometer (Iheanetu et al., 2019; Asad et al., 2021; Amiri et al., 2024), or by scanning a bright celestial calibrator for a single dish telescope (e.g. Jiang et al., 2020). Besides, the Tianlai experiment has measured the primary beam via unmanned aerial vehicle (Zhang et al., 2021) and CHIME has measured the primary beam via the Sun transition (Amiri et al., 2022).
In this work, we estimate the Stokes I primary beam patterns of the FAST telescope that combine the effects of its reflector and L-band feed array. The FAST L-band observation uses a 19-feed array mounted in the focus cabin suspended on the cables about the reflector. The reflector consists of about 4 40044004\,4004 400 active reflector panels, which form a $300$ m paraboloid at each pointing direction. Such a complicated reflector and feed arrangement heavily increases beam pattern instabilities, resulting in a direction-dependent beam pattern (Dong & Han, 2013). In this work, we developed a stacking-based beam pattern construction method employing Hi IM drift scan observation data itself. Because the beam shape is measured in the same direction as the Hi IM survey observation, we may minimize the direction dependency impact and achieve the beam pattern that is close to the actual case during the observation.
@@ 2 Data and preprocessing
Table 1: The observational data used in this work. In these observations, the low-level noise diode is adopted and the feed array is rotated to a fixed angle of $23.4^{\circ}$ during each observation to have an optimized sky coverage (Li et al., 2018).
@@
    The FATHOMER observations are carried out for $4$ hours drift-scan every night with $1$ second integration time. A low-level noise diode signal is injected for $1\rm s$ in every $8\,{\rm s}$.
@@
    The CRAFTS observations are carried out for $5$ hours drift-scan every nights with $0.2$ second integration time. A low-level noise diode signal is injected with high-cadence, injected for $81.92\mu{\rm s}$ in every $198.608\,{\rm\mu s}$..
@@
    The field center, encoded with ‘HHMM+ddmm’, indicates the center of each drift scan survey field, where ‘HHMM’ represents the R.A. and ‘+ddmm’ represents the declination, respectively.
@@ 2.1 FATHOMER
We employ the drift scan observation data from the FAst neuTral HydrOgen intensity Mapping ExpeRiment (FATHOMER; Li et al., 2023). In this analysis, $18\times 4$ hr drift scan observation data with integration time of $1$ s per timestamp are used, which targets the area that covers the R.A. range from 9 to 13 hr, overlapping with the Northern Galactic Cap (NCP) area of the Sloan Digital Sky Survey (SDSS; Reid et al., 2016). To achieve in-time calibration, a noise diode signal with known power is injected for $1\,{\rm s}$ in every $8\,{\rm s}$. We adopt a low-level noise diode  [^6] The FAST flux calibration uses a noise diode with two power levels: $\sim 1\,{\rm K}$ (low-level) and $\sim 10\,{\rm K}$ (high-level). For both FATHOMER and CRAFTS drift scan observations, the low-level noise diode injection is used. during the observations. The drift scan time-ordered data are preprocessed via the analysis pipeline developed in Li et al. (2023). Following the pipeline, the full frequency band is first split into three sub-bands, i.e. the low-frequency band 10501050105010501150115011501150 MHz, the mid-frequency band 11501150115011501250125012501250 MHz, and the high-frequency band 12501250125012501450145014501450 MHz. In this work, we use the high-frequency band data as it is less contaminated by RFI. The high-frequency band also covers the continuum spectrum range of the celestial point source catalog, which is used as a sky model for beam pattern construction. Extending the lower frequency band requires a better knowledge of the source spectrum model. The selected data are then RFI flagged and flux calibrated before proceeding to the analysis of this work.
@@ 2.2 CRAFTS
The Commensal Radio Astronomy FasT Survey (CRAFTS; Li et al., 2018) is designed to cover the FAST sky between declination of $-14^{\circ}$ and $+66^{\circ}$ and proposed to carry out multiple surveys simultaneously in a single drift scan observation, including Galactic Hi, extra-galactic Hi, pulsar search, and transient surveys. This analysis uses $14\times 5$ hr CRAFTS drift scan observation data, which targets the area that covers the R.A. range from $1$ to $6$ hr and the declination range from $27^{\circ}15^{\prime}$ to $32^{\circ}19^{\prime}$. CRAFTS drift scan observation data have the high-cadence noise diode signal injection777 To avoid data contamination in Fourier space for pulsar search, the noise injection period for CRAFTS is set to $198.608\mu{\rm s}$ with $41\%$ duty cycle, i.e. with $81.92\mu{\rm s}$ noise diode on in every injection period (Yang et al., 2024). and the noise diode temperature was pre-calibrated (Jiang et al., 2020). The pre-pipeline of CRAFTS data analysis can be found in Yang et al. (2024). The integration time for CRAFTS is $0.2$ s per timestamp. We reduce the integration time to $1$ s after the calibration processing. We adopt the same frequency partition strategy and time-ordered data analysis pipeline as the FATHOMS analysis. All the FAST drift scan data used in this work are summarized in Table 1.
@@ 2.3 NVSS
The FATHOMER and CRAFTS survey fields are both fully covered by the NVSS field. Both the FATHOMER and CRAFTS surveys adopt the drift scan observation mode, which is scanning at a constant declination within an R.A. range. During the drift scans, the rotation angle of the 19-feed array is fixed at $23.4^{\circ}$ to obtain the optimized field coverage (Li et al., 2018). Therefore, the same sources might be observed multiple times by different feeds. In addition, we also selected the sources with their declination within $10^{\prime}$ to the scanning declination to recover a full pattern beyond the size of the main beam. We choose a total area of $\sim 138\,{\rm\deg}^{2}$ and $\sim 385\,{\rm\deg^{2}}$ NVSS field for FATHOMER and CRAFTS, respectively. The total number of sources within the survey areas are 7 46374637\ 4637 463 and 21 0152101521\ 01521 015 for FATHOMER and CRAFTS, respectively. However, not all the sources are used in the beam pattern construction. Some of the sources are dropped because they are either poorly observed or highly contaminated by RFI or nearby sources. The detailed source selection criteria are presented in Section 3.3.
@@ 3.1 Beam stacking
where $\bm{d}$ and $\bm{n}$ are the time-ordered data vector and the noise vector, with the vector element representing the flux measurements and the corresponding noise at each time stamp, respectively. The multiplication of $\bm{M}\bm{b}$ represents the convolution of the sky model and the beam kernel. $\bm{b}=\{\bm{B}_{1}^{\rm T},\bm{B}_{2}^{\rm T},\cdots,\bm{B}_{q}^{\rm T}\}^{\rm T}$ is a column vector by stacking the columns of beam matrix $\bm{B}$ with size $(q,q)$, where $\bm{B}_{i}$ represents the $i$-th column of the beam matrix. Matrix $\bm{M}$ is a special matrix of size $(p,q^{2})$, where $p$ is the number of pixels within the survey area. The matrix $\bm{M}$ captures the convolutional spatial relationships between the pixels of the sky and the beam kernel matrix $\bm{B}$, i.e. each row of the matrix $\bm{M}$ is contracted by stacking the pixels covered by the beam kernel at each pointing direction. The matrix $\bm{P}$ is the projection matrix that maps the beam-convolved sky pixels to the time stamps.
@@
where $\hat{b}_{i,\cdots,j}$ represents a vector composed with the $i$-th to $j$-th elements of $\hat{\bm{b}}$. We use a beam matrix with the shape of $100\times 100$ and resolution of $12^{\prime\prime}$, covering $20^{\prime}$ in both dimensions, which is large enough to include the first few side lobes.
We adopt the time-ordered data (TOD) pre-processed in the previous work of Li et al. (2023) and Yang et al. (2024) for the FATHOMER and CRAFTS data, respectively. The data process is performed on the two linear polarization auto-power, referred to as $d_{\rm XX}$ and $d_{\rm YY}$, which is then combined to form the Stokes I components,
$d_{\rm I}=(d_{\rm XX}+d_{\rm YY})/2.$ (5)
It is worth noticing that, with the TOD pre-processing, the data is pre-whitened by subtracting the frequency-averaged temporal baseline.
@@ 3.2 The sky model
$m_{p}=\sum_{i\in\mathcal{I}}S_{i},\quad\mathcal{I}=\{i\mid|\alpha_{i}-\alpha_{p}|<\frac{\delta}{2},|\beta_{i}-\beta_{p}|<\frac{\delta}{2}\}$ (6)
where $(\alpha_{i},\beta_{i})$ represents the coordinate of the $i$-th sources with flux density of $S_{i}$ and $m_{p}$ is the total flux density of the pixel at $(\alpha_{p},\beta_{p})$ with pixel size of $\delta$. In this work, the pixel size of the map is set to match the beam model kernel, i.e. $\delta=12^{\prime\prime}$. The pixel size in this work is sufficiently small relative to the FAST beam width of $\sim 3^{\prime}$, ensuring minimal position accuracy loss due to pixelation.